Equivalent Units of Production and Related Costs

The charges to Work in Process—Assembly Department for a period, together with information concerning production, are as follows. All direct materials are placed in process at the beginning of production.

Work in Process—Assembly Department
Bal., 1,600 units, 35% completed 17,440 To Finished Goods, 29,600 units ?
Direct materials, 29,000 units @ $9.50 275,500
Direct labor 84,600
Factory overhead 39,258
Bal. ? units, 45% completed ?

Determine the following:

a. The number of units in work in process inventory at the end of the period.
units

b. Equivalent units of production for direct materials and conversion. If an amount is zero or a blank, enter in "0".

Work in Process-Assembly Department
Equivalent Units of Production for Direct Materials and Conversion Costs
Whole Units Equivalent Units
Direct Materials
Equivalent Units
Conversion
Inventory in process, beginning
Started and completed
Transferred to finished goods
Inventory in process, ending
Total units

c. Costs per equivalent unit for direct materials and conversion.

Costs Per Equivalent Unit
Direct Materials $
Conversion $

d. Cost of the units started and completed during the period.
$
Answer #1

a) Ending work in process unit = 1600+29000-29600 = 1000 Units

b. Equivalent units of production for direct materials and conversion. If an amount is zero or a blank, enter in "0".

Work in Process-Assembly Department
Equivalent Units of Production for Direct Materials and Conversion Costs
Whole Units Equivalent Units
Direct Materials
Equivalent Units
Conversion
Inventory in process, beginning 1600 0 1600*65% = 1040
Started and completed 28000 28000 28000
Transferred to finished goods 29600 28000 29040
Inventory in process, ending 1000 1000 1000*45% = 450
Total units 30600 29000 29490

c. Costs per equivalent unit for direct materials and conversion.

Direct material = 275500/29000 = 9.50

Conversion = (84600+39258)/29490 = 4.20

d) Cost of unit started and completed = (9.50+4.20)*28000 = $383600